ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60F15 Klesov Oleg, National Technical University, Kiev, Ukraine Fazekas, Istvan, Kossuth Lajos University, Debrecen, Hungary A general form of the strong law of large numbers We give conditions for maximal inequalities for sums of random variables. These inequalities allow one to obtain strong law of large numbers without any assumption on dependence structure of random variables. We present some applications of results obtained.